OpenStudy (anonymous):

Find the arc length of a sector that is 125° and has a radius of 20 in.

OpenStudy (anonymous):

I think i know how to do it but im not sure is it 3.14 x 20 x 20 = 1256 Arc length = 125/360 x 1256 = 436.11

OpenStudy (anonymous):

 Arc length is proportioanl to the angle in radians. \[L={\theta\over360}\times2\pi r\]

OpenStudy (anonymous):

So im right? with 436.11?

OpenStudy (anonymous):

you calculated the Area of the sector. by usgin \(\pi r^2\).

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Area of sector: \[A={\theta\over360}\times \pi r^2\] Length of arc: \[L={\theta\over360}\times2\pi r\]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

\[arclength=\theta \times r\] where r is the radius and theta is angle in radians

OpenStudy (anonymous):

okay yes i got it! thanks :)

OpenStudy (anonymous):

yer welcome.